OpenStudy (highschoolmom2010):

Leon’s bicycle wheels have a circumference of 2 m. What is his linear speed when the wheels rotate at 1 revolution per second?

rvc (rvc):

circumference = 2m v = r \(\omega\)

rvc (rvc):

\(\omega\) = \(\Large\rm\frac{2 \pi}{T}\)

OpenStudy (highschoolmom2010):

what was omega in there for

rvc (rvc):

the formula for velocity is : Velocity = Radius X Omega Velocity = \(\rm\Large\frac{R ~X ~2 \Pi}{T}\)

rvc (rvc):

2 pi R =circumference

OpenStudy (highschoolmom2010):

so im not getting this ive looked at it all morning :(

rvc (rvc):

Velocity =\(\rm\large\frac{circumference}{Time}\)

rvc (rvc):

@highschoolmom2010

rvc (rvc):

Velocity = \(\rm\large\frac{Circumference}{time})\

rvc (rvc):

Velocity = \(\rm\large\frac{circumference}{time})\

OpenStudy (irishboy123):